A bond pays a coupon (or interest) rate of 5 percent each year for five years, with a future (face) value of $200. If the bond were sold today, what would be the present value of the bond?

Multiple Choice

  • $145

  • $157

  • $200

  • $150
Answer #1

Future value (FV) = 200

Interest rate (i) = 5% = 0.05

PV = FV/(1 +i)5

= 200/(1 + 0.05)5

= 200/(1.05)5

= 200/(1.2762815625)

= 156.7052339

so PV of bond is $157